Mary is twice as old as Ann was when Mary was as old as Ann now is. Ann is 57. Mary is how old?
Answer by AnlytcPhil(1116)

>>Mary is twice as old as Ann was when Mary was as old as Ann now is.
Ann is 57. Mary is how old?<<
Let M = Mary's age now
Let x = the number of years ago we are talking about.
Mary's age now = M
Ann's age now = 57
Mary's age then = M-x
Ann's age then = 57-x

>>...Mary [now] is twice as old as Ann was [then]...<<
Translation: M = 2(57-x)

>>...Mary [then] was as old as Ann now is...<<
Translation: M-x = 57
So we have this system of equations:
M = 2(57-x)
M-x = 57
Solve the second equation for x
x = M - 57
Substitute M-57 for x in the first equation
M = 2[57-(M-57)]
M = 2(57-M+57)
M = 2(114-M)
M = 228 - 2M
3M = 228
M = 228/3 = 76
That's the answer, Mary is now 76.
To check solve for x
x = M - 57 = 76 - 57 = 19 years ago
Mary's age now = M = 76
Ann's age now = 57
Mary's age then = M-x = 76-19 = 57
Ann's age then = 57-x = 57-19 = 38
So back when Mary was 57, Ann was 38, and Mary, now 76, is
indeed twice Ann's age back then, 38.
Edwin